OpenStudy (anonymous):

locate absolute extreme of h(t)=(t)/(t-2) on closed interval [3,5]

OpenStudy (tkhunny):

Have you considered the 1st derivative? That will help you on your way, but the derivative doesn't know that chopped it off at 3 and 5.

OpenStudy (anonymous):

can yu walk me through the steps?

OpenStudy (tkhunny):

1st Derivative is all yours. Let's see what you get.

OpenStudy (anonymous):

\[\frac{ 2 }{ (t-2) ^2}\]

OpenStudy (anonymous):

from what i know of i tink i set that to 0 and find t right?

OpenStudy (tkhunny):

First, you get the derivative right. \(h(t) = \dfrac{t}{t-2}\) \(h'(t) = \dfrac{(t-2)(1) - t(1)}{(t-2)^{2}} = \dfrac{t - 2 - t}{(t-2)^{2}} = \dfrac{-2}{(t-2)^{2}}\) You should see the value of showing your work. Just pulling things out of the sky is not very useful.

OpenStudy (anonymous):

oh i forgot the negative

OpenStudy (anonymous):

and now i set it to 0?